# Optimal Rates in Continual Linear Regression via Increasing Regularization

### v1

@@ 5 Related work
Recent theoretical work on continual learning has studied the explicitly regularized scheme (Scheme 1) in continual linear regression settings [27, 47, 28], with several key differences from our work. Like we do, these papers focused on settings where labels stem from an underlying linear model. However, they analyzed the generalization loss given noisy data, while we analyze the training loss given noiseless data. Theirs may sound like a “stronger”, more permissive setup, but comes at the price of a very restrictive assumption: the expected task covariances $\mathbb{E}\mathbf{X}_{1}^{\top}\mathbf{X}_{1},\dots,\mathbb{E}\mathbf{X}_{M}^{\top}\mathbf{X}_{M}$ are assumed to commute. This commutativity removes forgetting due to misaligned feature subspaces across tasks, leaving noise as the sole culprit behind any degradation.
@@
Li et al. [27, 28] focused exclusively on sequences of $M=2$ tasks. Li et al. [27] derived risk bounds for isotropic regularization (Scheme 1) and highlight a trade-off between forgetting and intransigence. Li et al. [28] demonstrated that, under additional restrictions on the data matrices, there is a trade-off between increased memory usage and the performance of regularized continual linear regression. In all these works, performance degradation is attributed solely to label noise. In contrast, we analyzed interference that arises even in the absence of noise. Accordingly, their focus lies in a complementary regime that does not capture the challenges we address.
@@
Two main theoretical works studied the finite budget setting (Scheme 2). Jung et al. [22] analyzed continual linear classification under cyclic and random orderings. For cyclic orderings, they provided convergence rate for the loss; and, for random orderings, they only proved asymptotic convergence. Moreover, classification settings can yield different results and conclusions compared to regression settings [see 12]. Zhao et al. [47] analyzed both regularized and budgeted continual linear regression schemes under restrictive assumptions, showing that a carefully constructed, task-dependent regularization matrix can force the iterates of the regularized scheme to match those of the budgeted one. This alignment, however, requires precise knowledge of task covariances and breaks under standard isotropic $\ell_{2}$ regularization. In contrast, our unified reduction of both schemes to IGD (Section 3) avoids this limitation entirely.
@@
Cai and Diakonikolas [5] analyzed the Incremental Proximal Method (IPM), corresponding to isotropic $\ell_{2}$ regularization, under cyclic orderings. They provided convergence rates for convex smooth or convex Lipschitz losses with bounded noise, but their guarantees only become meaningful after multiple full sweeps (or epochs) over the task sequence. In contrast, we analyze the random orderings and establish nontrivial—and even optimal—guarantees without requiring repeated passes. See Section 6 for a comparison with our regularization schedules.
@@ 6 Discussion
Our work reduces regularized continual linear regression—whether using explicit or implicit regularization—to incremental gradient descent on smooth surrogate losses. This reduction enables last-iterate SGD analysis under random task orderings and yields significantly improved convergence rates. Specifically, we show that a suitable fixed regularization strength achieves a near-optimal rate of $\mathcal{O}({\log k}/{k})$, while a carefully chosen increasing strength schedule achieves, for the first time, an $\mathcal{O}({1}/{k})$ rate, matching the known lower bound. See Figures 1 and 1 for a summary of the reductions and improved rates.
@@
In Section 4.2, we derived an optimal regularization schedule in which the regularization strength increases with each task. This implies that the parameters change progressively less over time. Interestingly, such an attenuation in “synaptic plasticity” is also observed in biological systems: the rate at which synapses grow or shrink in response to sensory stimulation [43] or motor learning [20] significantly decreases over time as the brain matures [34].
In continual learning, many papers practically set a fixed regularization coefficient $\lambda$ through simple hyperparameter tuning. However, non-isotropic weighting schemes often encode an implicit scale in the weighting matrices they compute. Methods such as EWC [24] and Path Integral [46] are particularly sensitive to $\lambda$, as their weighting matrices tend to have low magnitude early in training and may increase over time [see 9]. This initially low regularization strength was considered problematic by some [e.g., 6] and was even canceled algorithmically, as it allows excessive plasticity in early tasks. Yet, one may argue that high plasticity is desirable in the beginning of long task sequences, where substantial expertise remains to be acquired. Our analysis in Theorem 4.6 supports this intuition, showing that in such cases, an increasing regularization schedule yields optimal upper bounds under random task orderings. See also the findings and discussion in Mirzadeh et al. [33] on the effects of a decaying step size, which—as noted in our Section 2—corresponds to an increasing regularization strength.
Analytically, Evron et al. [12] showed that in continual linear models for binary classification, increasing the regularization coefficient can be harmful to convergence guarantees (see their Example 3). However, their analysis applies only to weakly regularized schemes (where $\lambda_{t}\to 0$ for all $t$), and the problematic schedule they presented increases the coefficient at a doubly-exponential rate—in contrast to our Theorem 4.6 which utilizes finite, and relatively large, coefficients that increase linearly. Under cyclic orderings over linear regression tasks, solved with explicit regularization (Scheme 1), the analysis of Cai and Diakonikolas [5] dictates a fixed coefficient $\lambda=2MR^{2}\sqrt{\ln({k}/{M})}$. In contrast, under random orderings, our fixed variant in Section 4.1 sets $\lambda=R^{2}(\ln k-1)$. While both choices grow at most logarithmically with $k$, theirs grows with the number of tasks $M$, making it less suitable for “single-epoch” settings—though effective in the multi-epoch regime that they studied.
@@
Throughout the paper, we assumed Scheme 1 uses isotropic $\ell_{2}$ regularization. Such regularization often performs competitively with weighted schemes in practice [31, 39]. The latter, widely used in the literature, typically rely on weight matrices derived from Fisher information, often approximated by their diagonal [24, 46, 2, 4]. Theoretically, using the full Fisher matrix from previous tasks requires $\mathcal{O}(d^{2})$ memory in the worst case, but guarantees zero seen-task loss (Definition 4.8)—that is, complete retention of past expertise (see Proposition 5.5 of Evron et al. [12] and Proposition 5 of Peng et al. [35]).
@@
Our Lemma 4.7, originally proved to leverage the reductions from continual learning to the incremental gradient descent method, also establishes a last-iterate convergence guarantee for a variant of SGD that may be of independent interest. By setting the surrogate functions equal to the original functions, this result yields an $\mathcal{O}(1/k)$ convergence guarantee for convex smooth optimization in the realizable regime, using a linear decay schedule [8]. To our knowledge, this is the first fast-rate guarantee for the last-iterate convergence of SGD in the realizable setting. It not only generalizes prior results specific to least-squares problems [42], but also improves the convergence rate from $O(\log T/T)$ to the optimal $O(1/T)$.
@@
While our analysis establishes optimal rates for realizable continual linear regression with regularization under random task orderings, several directions remain open. First, empirical validation on standard continual learning benchmarks would test the practical impact of our regularization strength choices. Second, extending our reduction-based analysis to simple nonlinear models could reveal whether similar regularization schedules yield optimal convergence in more expressive settings.
Perhaps more fundamentally, it remains an open question whether comparable bounds can be achieved without regularization—that is, whether the gap between the upper bound of Evron et al. [13] and the lower bound—both shown in Table 1—can be closed. Clarifying this would shed light on whether regularization is essential for achieving optimal rates, or if it simply serves a role as a convenient analytical device. Addressing these questions may help close the gap between theoretical guarantees and practical continual learning systems.
@@
Appendix A Proofs of lower bounds
Theorem A.1.

### v2

@@ 5 Related work
Recent theoretical work on continual learning has studied the explicitly regularized scheme (Scheme˜1) in continual linear regression settings [28, 48, 29], with several key differences from our work. Like we do, these papers focused on settings where labels stem from an underlying linear model. However, they analyzed the generalization loss given noisy data, while we analyze the training loss given noiseless data. Theirs may sound like a “stronger”, more permissive setup, but comes at the price of a very restrictive assumption: the expected task covariances $\mathbb{E}\mathbf{X}_{1}^{\top}\mathbf{X}_{1},\dots,\mathbb{E}\mathbf{X}_{M}^{\top}\mathbf{X}_{M}$ are assumed to commute. This commutativity removes forgetting due to misaligned feature subspaces across tasks, leaving noise as the sole culprit behind any degradation.
@@
Li et al. [28, 29] focused exclusively on sequences of $M=2$ tasks. Li et al. [28] derived risk bounds for isotropic regularization (Scheme˜1) and highlight a trade-off between forgetting and intransigence. Li et al. [29] demonstrated that, under additional restrictions on the data matrices, there is a trade-off between increased memory usage and the performance of regularized continual linear regression. In all these works, performance degradation is attributed solely to label noise. In contrast, we analyzed interference that arises even in the absence of noise. Accordingly, their focus lies in a complementary regime that does not capture the challenges we address.
@@
Two main theoretical works studied the finite budget setting (Scheme˜2). Jung et al. [23] analyzed continual linear classification under cyclic and random orderings. For cyclic orderings, they provided convergence rate for the loss; and, for random orderings, they only proved asymptotic convergence. Moreover, classification settings can yield different results and conclusions compared to regression settings [see 13]. Zhao et al. [48] analyzed both regularized and budgeted continual linear regression schemes under restrictive assumptions, showing that a carefully constructed, task-dependent regularization matrix can force the iterates of the regularized scheme to match those of the budgeted one. This alignment, however, requires precise knowledge of task covariances and breaks under standard isotropic $\ell_{2}$ regularization. In contrast, our unified reduction of both schemes to IGD (Section˜3) avoids this limitation entirely.
@@
Cai and Diakonikolas [5] analyzed the Incremental Proximal Method (IPM), corresponding to isotropic $\ell_{2}$ regularization, under cyclic orderings. They provided convergence rates for convex smooth or convex Lipschitz losses with bounded noise, but their guarantees only become meaningful after multiple full sweeps (or epochs) over the task sequence. In contrast, we analyze the random orderings and establish nontrivial—and even optimal—guarantees without requiring repeated passes. See Section˜6 for a comparison with our regularization schedules.
@@ 6 Discussion
Our work reduces regularized continual linear regression—whether using explicit or implicit regularization—to incremental gradient descent on smooth surrogate losses. This reduction enables last-iterate SGD analysis under random task orderings and yields significantly improved convergence rates. Specifically, we show that a suitable fixed regularization strength achieves a near-optimal rate of $\mathcal{O}({\log k}/{k})$, while a carefully chosen increasing strength schedule achieves, for the first time, an $\mathcal{O}({1}/{k})$ rate, matching the known lower bound. See Figures˜1 and 1 for a summary of the reductions and improved rates.
@@
In Section˜4.2, we derived an optimal regularization schedule in which the regularization strength increases with each task. This implies that the parameters change progressively less over time. Interestingly, such an attenuation in “synaptic plasticity” is also observed in biological systems: the rate at which synapses grow or shrink in response to sensory stimulation [44] or motor learning [21] significantly decreases over time as the brain matures [35].
In continual learning, many papers practically set a fixed regularization coefficient $\lambda$ through simple hyperparameter tuning. However, non-isotropic weighting schemes often encode an implicit scale in the weighting matrices they compute. Methods such as EWC [25] and Path Integral [47] are particularly sensitive to $\lambda$, as their weighting matrices tend to have low magnitude early in training and may increase over time [see 10]. This initially low regularization strength was considered problematic by some [e.g., 6] and was even canceled algorithmically, as it allows excessive plasticity in early tasks. Yet, one may argue that high plasticity is desirable in the beginning of long task sequences, where substantial expertise remains to be acquired. Our analysis in Theorem˜4.6 supports this intuition, showing that in such cases, an increasing regularization schedule yields optimal upper bounds under random task orderings. See also the findings and discussion in Mirzadeh et al. [34] on the effects of a decaying step size, which—as noted in our Section˜2—corresponds to an increasing regularization strength.
Analytically, Evron et al. [13] showed that in continual linear models for binary classification, increasing the regularization coefficient can be harmful to convergence guarantees (see their Example 3). However, their analysis applies only to weakly regularized schemes (where $\lambda_{t}\to 0$ for all $t$), and the problematic schedule they presented increases the coefficient at a doubly-exponential rate—in contrast to our Theorem˜4.6 which utilizes finite, and relatively large, coefficients that increase linearly. Under cyclic orderings over linear regression tasks, solved with explicit regularization (Scheme˜1), the analysis of Cai and Diakonikolas [5] dictates a fixed coefficient $\lambda=2MR^{2}\sqrt{\ln({k}/{M})}$. In contrast, under random orderings, our fixed variant in Section˜4.1 sets $\lambda=R^{2}(\ln k-1)$. While both choices grow at most logarithmically with $k$, theirs grows with the number of tasks $M$, making it less suitable for “single-epoch” settings—though effective in the multi-epoch regime that they studied.
To complement our analysis, App.˜A empirically examines a simple two-task setup trained under the explicit-regularization scheme with random task ordering. Interestingly, in this non-worst-case setting, the optimal fixed regularization, obtained by minimizing the expected loss after $k$ steps, exhibits an approximately linear dependence on $k$, in contrast to the logarithmic scaling we predicted for the worst case.
@@
Throughout the paper, we assumed Scheme˜1 uses isotropic $\ell_{2}$ regularization. Such regularization often performs competitively with weighted schemes in practice [32, 40]. The latter, widely used in the literature, typically rely on weight matrices derived from Fisher information, often approximated by their diagonal [25, 47, 2, 4]. Theoretically, using the full Fisher matrix from previous tasks requires $\mathcal{O}(d^{2})$ memory in the worst case, but guarantees zero seen-task loss (Definition˜4.8)—that is, complete retention of past expertise (see Proposition 5.5 of Evron et al. [13] and Proposition 5 of Peng et al. [36]). Closed-form alternatives include Recursive Least Squares (8, Chapter 12.2 in) and its block variant BRMP [49], which likewise maintain an explicit second-order memory to prevent forgetting. More generally, the rank of the regularization weight matrix can offer a trade-off between memory and forgetting [29].
@@
Our Lemma˜4.7, originally proved to leverage the reductions from continual learning to the incremental gradient descent method, also establishes a last-iterate convergence guarantee for a variant of SGD that may be of independent interest. By setting the surrogate functions equal to the original functions, this result yields an $\mathcal{O}(1/k)$ convergence guarantee for convex smooth optimization in the realizable regime, using a linear decay schedule [9]. To our knowledge, this is the first fast-rate guarantee for the last-iterate convergence of SGD in the realizable setting. It not only generalizes prior results specific to least-squares problems [43], but also improves the convergence rate from $O(\log T/T)$ to the optimal $O(1/T)$.
@@
While we establish optimal rates for realizable continual linear regression with regularization under random task orderings, several directions remain open. First, empirical validation on standard continual learning benchmarks would test the practical impact of our regularization strength schedules. Second, extending our reduction-based analysis to simple nonlinear models could reveal whether similar schedules yield optimal convergence in more expressive settings.
Perhaps more fundamentally, it remains an open question whether comparable bounds can be achieved without regularization—that is, whether the gap between the upper bound of Evron et al. [14] and the lower bound—both shown in Table˜1—can be closed. Clarifying this would shed light on whether regularization is essential for achieving optimal rates, or if it simply serves a role as a convenient analytical device. Addressing these questions may help close the gap between theoretical guarantees and practical continual learning systems. Finally, fully understanding the role of regularization in typical (non-worst-case) settings remains an open question for future work.
@@
Appendix A Empirical Illustration
In this section, we illustrate the dependence of the optimal fixed regularization coefficient $\lambda$ on the horizon $k$ and task angle $\theta$ in a simple two-task setup. Two single sample tasks are defined as $\mathbf{x}_{1}=(1,0)$ and $\mathbf{x}_{2}=(\cos\theta,\sin\theta)$, trained under the explicit–regularization scheme (Scheme˜1) with a random with–replacement ordering (Definition˜4.2). For each $(k,\theta)$ we numerically determine the fixed coefficient $\lambda^{\star}(k;\theta)$ that minimizes the expected loss after $k$ steps. As shown in Figure˜2, the optimal regularization grows roughly linearly with $k$, and its magnitude increases with $\theta$–indicating that longer horizons and more misaligned tasks benefit from stronger regularization.
Figure 2: Optimal fixed regularization grows with horizon and task angle. Each curve shows $\lambda^{\star}(k;\theta)$ obtained by minimizing the expected loss after $k$ steps of the explicit-regularization scheme with a constant $\lambda$. We observe an approximately linear growth in $k$ and higher optimal regularization for larger $\theta$.
Discussion.
The qualitative behavior in Figure˜2 aligns with the general principles established in the paper, but the quantitative trend differs. In Section˜4, we prove an optimal time–varying regularization schedule and a near-optimal fixed coefficient for the worst-case analysis over arbitrary task distributions. Here, by contrast, we examine a simple two-task setting, where the empirically optimal fixed regularization $\lambda^{\star}(k;\theta)$ grows approximately linearly with $k$, rather than logarithmically as predicted by Corollary˜4.4. A plausible explanation is that the worst-case distribution inducing the $\log(k)$ scaling is more complex than the two-task setup considered here, so optimizing for this simple distribution yields an inherently different solution from the true worst-case behavior across all distributions.
Computation.
We compute the expected loss in closed form using an operator formulation similar to that of Evron et al. [14], originally developed to establish their fast dimensionality-dependent rate. At each step of Scheme˜1,
${\mathbf{w}}_{t}=(\mathbf{X}_{\tau_{t}}^{\top}\mathbf{X}_{\tau_{t}}+\lambda\mathbf{I})^{-1}(\mathbf{X}_{\tau_{t}}^{\top}\mathbf{y}_{\tau_{t}}+\lambda{\mathbf{w}}_{t-1}),\quad\lambda>0.$
Under realizability ($\mathbf{y}_{m}=\mathbf{X}_{m}{\mathbf{w}}_{\star}$), defining ${\mathbf{z}}_{t}={\mathbf{w}}_{t}-{\mathbf{w}}_{\star}$ gives
${\mathbf{z}}_{t}$ $=(\mathbf{X}_{\tau_{t}}^{\top}\mathbf{X}_{\tau_{t}}+\lambda\mathbf{I})^{-1}(\mathbf{X}_{\tau_{t}}^{\top}\mathbf{y}_{\tau_{t}}+\lambda{\mathbf{w}}_{t-1})-{\mathbf{w}}_{\star}$
$=(\mathbf{X}_{\tau_{t}}^{\top}\mathbf{X}_{\tau_{t}}+\lambda\mathbf{I})^{-1}(\mathbf{X}_{\tau_{t}}^{\top}\mathbf{X}_{\tau_{t}}{\mathbf{w}}_{\star}+\lambda{\mathbf{w}}_{t-1})-{\mathbf{w}}_{\star}$
$=(\mathbf{X}_{\tau_{t}}^{\top}\mathbf{X}_{\tau_{t}}+\lambda\mathbf{I})^{-1}\big((\mathbf{X}_{\tau_{t}}^{\top}\mathbf{X}_{\tau_{t}}+\lambda\mathbf{I}){\mathbf{w}}_{\star}+\lambda({\mathbf{w}}_{t-1}-{\mathbf{w}}_{\star})\big)-{\mathbf{w}}_{\star}$
$={\mathbf{w}}_{\star}+\lambda(\mathbf{X}_{\tau_{t}}^{\top}\mathbf{X}_{\tau_{t}}+\lambda\mathbf{I})^{-1}({\mathbf{w}}_{t-1}-{\mathbf{w}}_{\star})-{\mathbf{w}}_{\star}$
$=\lambda(\mathbf{X}_{\tau_{t}}^{\top}\mathbf{X}_{\tau_{t}}+\lambda\mathbf{I})^{-1}{\mathbf{z}}_{t-1}.$
Since $\tau_{t}$ is sampled independently at each step, taking expectations over the randomness of $\tau$ gives
$\mathbb{E}[{\mathbf{z}}_{t}{\mathbf{z}}_{t}^{\top}]=\mathbb{E}\!\left[\lambda(\mathbf{X}_{\tau_{t}}^{\top}\mathbf{X}_{\tau_{t}}+\lambda\mathbf{I})^{-1}\,\mathbb{E}[{\mathbf{z}}_{t-1}{\mathbf{z}}_{t-1}^{\top}]\,\lambda(\mathbf{X}_{\tau_{t}}^{\top}\mathbf{X}_{\tau_{t}}+\lambda\mathbf{I})^{-1}\right].$
Starting from ${\mathbf{w}}_{0}=\mathbf{0}$, the initial error is ${\mathbf{z}}_{0}=-{\mathbf{w}}_{\star}$, so $\mathbb{E}[{\mathbf{z}}_{0}{\mathbf{z}}_{0}^{\top}]={\mathbf{w}}_{\star}{\mathbf{w}}_{\star}^{\top}$. Define the linear operator
$\mathcal{Q}_{\lambda}:\ \mathbb{R}^{d\times d}\to\mathbb{R}^{d\times d},\qquad\mathcal{Q}_{\lambda}[\mathbf{A}]=\mathbb{E}_{\tau}\!\left[\lambda(\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}+\lambda\mathbf{I})^{-1}\,\mathbf{A}\,\lambda(\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}+\lambda\mathbf{I})^{-1}\right].$
By independence of $\tau_{t}$ and ${\mathbf{z}}_{t-1}$,
$\mathbb{E}[{\mathbf{z}}_{t}{\mathbf{z}}_{t}^{\top}]=\mathcal{Q}_{\lambda}\!\left[\mathbb{E}[{\mathbf{z}}_{t-1}{\mathbf{z}}_{t-1}^{\top}]\right],\qquad\mathbb{E}[{\mathbf{z}}_{k}{\mathbf{z}}_{k}^{\top}]=\mathcal{Q}_{\lambda}^{\,k}\!\left[{\mathbf{w}}_{\star}{\mathbf{w}}_{\star}^{\top}\right].$
By Definition˜2.1, the population loss after $k$ steps is the expected training loss across tasks:
$\mathbb{E}[\mathcal{L}({\mathbf{w}}_{k})]=\frac{1}{2}\mathbb{E}\!\big[\|\mathbf{X}_{\tau}{\mathbf{w}}_{k}-\mathbf{y}_{\tau}\|^{2}\big].$
Under realizability ($\mathbf{y}_{\tau}=\mathbf{X}_{\tau}{\mathbf{w}}_{\star}$), this equals
$\mathbb{E}[\mathcal{L}({\mathbf{w}}_{k})]=\frac{1}{2}\mathbb{E}\!\big[\|\mathbf{X}_{\tau}({\mathbf{w}}_{k}-{\mathbf{w}}_{\star})\|^{2}\big]=\frac{1}{2}\mathbb{E}\!\big[{\mathbf{z}}_{k}^{\top}\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}{\mathbf{z}}_{k}\big].$
Taking the expectation inside the trace gives
$\mathbb{E}[\mathcal{L}({\mathbf{w}}_{k})]=\tfrac{1}{2}\mathrm{Tr}\!\Big(\mathbb{E}[\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}]\;\mathbb{E}[{\mathbf{z}}_{k}{\mathbf{z}}_{k}^{\top}]\Big)=\tfrac{1}{2}\mathrm{Tr}\!\Big(\mathbb{E}[\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}]\;\mathcal{Q}_{\lambda}^{\,k}[{\mathbf{w}}_{\star}{\mathbf{w}}_{\star}^{\top}]\Big).$
To move $\mathcal{Q}_{\lambda}^{\,k}$ from ${\mathbf{w}}_{\star}{\mathbf{w}}_{\star}^{\top}$ onto $\mathbb{E}[\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}]$, expand $\mathcal{Q}_{\lambda}^{\,k}$ along an i.i.d. sequence $\tau_{1},\ldots,\tau_{k}$ and apply linearity of expectation and the cyclic property of the trace:
$\mathrm{Tr}\!\Big(\mathbb{E}[\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}]\;\mathcal{Q}_{\lambda}^{\,k}[{\mathbf{w}}_{\star}{\mathbf{w}}_{\star}^{\top}]\Big)$ $=\mathbb{E}_{\tau_{1},\ldots,\tau_{k}}\!\Big[\mathrm{Tr}\!\big(\mathbb{E}[\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}]\,\bm{\Lambda}_{\tau_{k}}\cdots\bm{\Lambda}_{\tau_{1}}\,{\mathbf{w}}_{\star}{\mathbf{w}}_{\star}^{\top}\,\bm{\Lambda}_{\tau_{1}}\cdots\bm{\Lambda}_{\tau_{k}}\big)\Big]$
$=\mathbb{E}_{\tau_{1},\ldots,\tau_{k}}\!\Big[\mathrm{Tr}\!\big(\bm{\Lambda}_{\tau_{1}}\cdots\bm{\Lambda}_{\tau_{k}}\,\mathbb{E}[\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}]\,\bm{\Lambda}_{\tau_{k}}\cdots\bm{\Lambda}_{\tau_{1}}\,{\mathbf{w}}_{\star}{\mathbf{w}}_{\star}^{\top}\big)\Big]$
$=\mathrm{Tr}\!\Big(\mathcal{Q}_{\lambda}^{\,k}\!\big[\mathbb{E}[\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}]\big]\;{\mathbf{w}}_{\star}{\mathbf{w}}_{\star}^{\top}\Big),$
where $\bm{\Lambda}_{\tau}:=\lambda(\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}+\lambda\mathbf{I})^{-1}$. Hence,
$\mathbb{E}[\mathcal{L}({\mathbf{w}}_{k})]=\tfrac{1}{2}\mathrm{Tr}\!\Big(\mathcal{Q}_{\lambda}^{\,k}\!\big[\mathbb{E}[\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}]\big]\;{\mathbf{w}}_{\star}{\mathbf{w}}_{\star}^{\top}\Big)=\tfrac{1}{2}\,{\mathbf{w}}_{\star}^{\top}\mathcal{Q}_{\lambda}^{\,k}\!\big[\mathbb{E}[\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}]\big]{\mathbf{w}}_{\star}.$
Taking the maximum over all unit teachers $\|{\mathbf{w}}_{\star}\|=1$ gives the worst–case expected loss
$L_{k}^{\mathrm{worst}}=\tfrac{1}{2}\,\lambda_{\max}\!\Big(\mathcal{Q}_{\lambda}^{\,k}\!\big[\mathbb{E}[\mathbf{X}_{\tau}^{\top}\mathbf{X}_{\tau}]\big]\Big).$
This formulation allows us to compute the expected worst–case loss for Scheme˜1 entirely in closed form. All stochastic effects of task ordering are captured analytically through expectations, eliminating the need for Monte-Carlo sampling or repeated randomized experiments.
Appendix B Proofs of lower bounds
Theorem B.1.